at what speed the mass of proton be doubled
1
Expert's answer
2020-03-27T09:07:45-0400

The mass of a proton at speed v is


"m=2m_0=\frac{m_0}{\sqrt{1-\frac{v^2}{c^2}}}"

"2=\frac{1}{\sqrt{1-\frac{v^2}{c^2}}}\to 1-\frac{v^2}{c^2}=0.25"

"v=\frac{\sqrt{3}}{2}c=2.598\cdot10^8\frac{m}{s}"